OpenStudy (anonymous):

how do u solve for X ? 5x+6=2x-5

mathslover (mathslover):

\[\large{5x+6=2x-5}\] \[\large{5x+6-2x=2x-5-2x}\]

mathslover (mathslover):

Subtracting 2x both sides

mathslover (mathslover):

\[\large{3x+6=-5}\] \[\large{3x=-5-6}\] \[\large{3x=-11}\]

OpenStudy (anonymous):

if 3x = -11 then x would equal -11/3 correct??